OpenStudy (anonymous):

What is the quotient of c^2-c-2/c^2-1 divided by c^2+3c +2/c^2-3c+2 in lowest terms? A. (C-2)/(c-1) B. (C+1)(c+2)/(c-1)^2 C. (C-2)^2/(c+1)(c+2) D. (C-2)/(c+1)

OpenStudy (zenmo):

hello

OpenStudy (anonymous):

Hi @Zenmo

OpenStudy (zenmo):

Apply the reciprocal identity for division \[\frac{ \frac{ a }{ b } }{\frac{ c }{ d } } = \frac{ a }{ b } \times \frac{ d }{ c }\] to your equation.

OpenStudy (anonymous):

So C^2-c-2/c^2-1 x c^2-3c+2/c2+3c+2 @Zenmo

OpenStudy (zenmo):

Yes. Now, we look for "stuffs" that can be factored. Do you know how to factor an equation like \[x^2+7x+10\]?

OpenStudy (anonymous):

I can if you give me list of options. If there isn't then no... Not that I remember of. @Zenmo

OpenStudy (zenmo):

\[x^2+7x+10\] Let us start with 10: what two numbers multiply into 10?

OpenStudy (zenmo):

List all the combinations

OpenStudy (anonymous):

5 x 2 1 x 10 @Zenmo

OpenStudy (zenmo):

Next, look at the following set of numbers that multiply to 10, which set adds to 7?

OpenStudy (anonymous):

5 & 7. @Zenmo

OpenStudy (zenmo):

5 x 7 = 35 5 + 7 = 12 These do not meet the requirement of two numbers adding to 7 while also multiplying to 10.

OpenStudy (anonymous):

So oops I meant 5 +2 sorry @zenmo

OpenStudy (zenmo):

Why 5 and 2?

OpenStudy (anonymous):

It adds up to 7 and multiples to 10. So the factoring would be (x+5)(x+2) right? @zenmo

OpenStudy (zenmo):

Yes, and you understood on how you factored it?

OpenStudy (anonymous):

Yes. @zenmo

OpenStudy (zenmo):

Now, let us go back to your equation, \[c^2-c-2\]. Look at the 2nd term of " -c " and the 3rd term " 2 "

OpenStudy (zenmo):

what is the number of the 2nd term?

OpenStudy (anonymous):

-1 @zenmo

OpenStudy (zenmo):

Yes, what is the number of the third term?

OpenStudy (zenmo):

Let me know if you are stuck.

OpenStudy (anonymous):

-2. @zenmo. Soo (x+1)(x-2) right?

OpenStudy (anonymous):

For the factoring.. & I just noticed you didn't even ask for the factors.. Sorry. @zenmo

OpenStudy (zenmo):

Yes

OpenStudy (zenmo):

\[\frac{ (x+1)(x-2) }{ c^2-1 } \times \frac{ c^2-3c+2 }{ c^2+3c+2 }\]

OpenStudy (zenmo):

Good so far?

OpenStudy (anonymous):

Yes. @zenmo

OpenStudy (zenmo):

Now, we look for other parts that can be factored. Let us finished the left side of the multiplication. What is the factor of \[c^2-1\]?

OpenStudy (zenmo):

Hint: \[a^2-b^2 = (a+b)(a-b)\] Identity

OpenStudy (anonymous):

(C+1)(c-1) @zenmo

OpenStudy (zenmo):

Yes, correct. Now, the right side also looks to be factorable. Let us start with the numerator. Can you factor c2-3c+2

OpenStudy (anonymous):

Yeah. (X-2)(X-1) right? @zenmo

OpenStudy (zenmo):

Yes, correct, now, there is only one remaining. Factor c2+3c+2

OpenStudy (anonymous):

(X+1)(X+2) @zenmo

OpenStudy (zenmo):

\[\frac{ (c+1)(c-2) }{ (c+1)(c-1) }\times \frac{ (c-2)(c-1) }{ (c+2)(c+1) }\] it is not factored based on the work we just did.

OpenStudy (zenmo):

Do you know how to cancel?

OpenStudy (zenmo):

it is now factored*

OpenStudy (anonymous):

I think I know how to cancel, but I probably don't.... @zenmo

OpenStudy (zenmo):